OpenStudy (anonymous):

square root of 100x^3y^4

OpenStudy (anonymous):

i presume you mean: \[\sqrt{100x ^{3}y ^{4}}=\sqrt{100}\sqrt{x^3}\sqrt{y^4}=10x ^{\frac{ 3 }{ 2 }}y^2\]